OpenStudy (qwertykingghlpo):

f(x) = 9x^3 + 2x^2 - 5x + 4 and g(x) =5x^3 - 7x + 4 ... What is f(x) - g(x)? Pls show step by step :)

OpenStudy (qwertykingghlpo):

@Mahoganie.Carson

OpenStudy (qwertykingghlpo):

@ILovePuppiesLol

OpenStudy (qwertykingghlpo):

@ganeshie8

OpenStudy (qwertykingghlpo):

@PillowFace750

OpenStudy (qwertykingghlpo):

@shamil98 (this is appropriate)

OpenStudy (liz_beauty45):

wow people are working at this time??

OpenStudy (qwertykingghlpo):

haha ye.

OpenStudy (usukidoll):

hey there. so the first step is to put our equation in f(x) - g(x) form since f(x) = 9x^3+2x^2-5x+4 and g(x) = 5x^3-7x+4 f(x) -g(x) must be 9x^3+2x^2-5x+4 -( 5x^3-7x+4) but that's just the first step next step is to distribute the negative on the right hand side of the equation third and final step is to combine like terms

OpenStudy (qwertykingghlpo):

mkay :P

OpenStudy (usukidoll):

second step distributing the - on the right hand side of the equation 9x^3+2x^2-5x+4 -( 5x^3-7x+4) 9x^3+2x^2-5x+4 -5x^3+7x-4

OpenStudy (usukidoll):

we can also rearrange the equation to make it easier to combine like terms so instead of 9x^3+2x^2-5x+4 -5x^3+7x-4 we have 9x^3-5x^3+2x^2-5x+7x+4-4 now just subtract like terms and we're done

OpenStudy (qwertykingghlpo):

@UsukiDoll You freakin rock! :P

OpenStudy (usukidoll):

thanks. can you handle the problem from here? or do you need further assistance?

OpenStudy (qwertykingghlpo):

um could you do this last part just to make sure I get it right? :)

OpenStudy (usukidoll):

sure 9x^3-5x^3+2x^2-5x+7x+4-4 4x^3+2x^2+2x

OpenStudy (qwertykingghlpo):

Thank you :)

OpenStudy (qwertykingghlpo):

is this in factored form? @UsukiDoll

Directrix (directrix):

No. Not in factored form. 4x^3+2x^2+2x = 2x * ( 2x^2 + x + 1) @qwertykingGHLPO

OpenStudy (usukidoll):

I can't believe I got this late. ugh os lag. anyway the 2x is taken out and you're left with what Directrix wrote

OpenStudy (qwertty123):

:D